# General spin models from noncollinear spin density functional theory and spin-cluster expansion

### v1

@@ II.4 Real spherical harmonics represented as polynomials of Cartesian coordinates
In torque-fitted regression formalism, we must evaluate gradients of the spherical harmonics with respect to the unit vector, cf. Eq. (29). Differentiation in spherical coordinates $(\theta,\phi)$ is numerically ill-conditioned near the poles ($\sin\theta\!\to\!0$). To avoid these coordinate singularities, we represent the spherical harmonics as polynomials in the Cartesian components of the unit vector, which yields well-behaved analytic derivatives via $\partial/\partial e_{x}$, $\partial/\partial e_{y}$, and $\partial/\partial e_{z}$. The gradient of the complex spherical harmonics with respect to the unit vector was already derived in Ref. [30]. Here, we present expressions for the gradients of the real spherical harmonics in their Cartesian-polynomial form. See also the Supplementary Information for details of the derivations.
@@
II.5 Sampling method based on the mean-field approximation
Constructing an SCE model requires sampling spin configurations that adequately span the magnetic states of interest. In general, building a model that reproduces energies across arbitrary magnetic states with high accuracy within a realistic computational budget will demand sophisticated strategies, which we defer to future work. In the applications presented below, we instead assume that the relevant magnetic states are known a priori. For this setting, we employ a practical mean-field-based sampling scheme for the Heisenberg model that efficiently generates spin configurations representative of the target states [11, 31, 32]. Within the mean-field Heisenberg model single-site orientational probability $P(\hat{\bm{e}}_{i})$ takes the von Mises–Fisher form,
$P(\hat{\bm{e}}_{i})=\cfrac{3m/\tau}{4\pi\sinh(3m/\tau)}\exp\left(\cfrac{3m\hat{\bm{e}}\cdot\hat{\bm{e}}_{i}}{\tau}\right),$ (46)
where $^{\bm{m}}=\langle\hat{\bm{e}}\rangle=m\hat{\bm{e}}$ is the site-independent thermal average of the spin and $\tau=T/T_{\rm c}^{\rm MFA}$ is a reduced temperature ($T$ is the temperature and $T_{\rm c}^{\rm MFA}$ is the mean-field critical temperature). The parameters $m$ and $\tau$ are self-consistently related; fixing one automatically determines the other. Therefore, by choosing $\tau$, one can sample $\hat{\bm{e}}_{i}$ from this distribution to generate spin configurations representative of the target magnetic disorder within the mean-field approximation. Please refer to Ref. [11] for the derivation and further details.
II.6 Application to B20-type ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$
@@
B20-type chiral magnetic compounds ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$ exhibit helical spin structures [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. From a theoretical standpoint, their spin–orbit coupling is moderate, so they serve as testbeds for developing DFT-based methods to evaluate the Dzyaloshinskii–Moriya interaction (DMI)  [44, 45, 46, 47]. Here, we apply the SCE method to the B20-type compounds and examine the applicability via the evaluation of exchange interaction, DMI, and helical spin period.
We used a $\times2$$\times$2 cubic supercell including 64 atoms with experimental crystallographic structural parameters of FeGe [34] and constructed an SCE model using spin configurations sampled from states close to the ferromagnetic state. Sampling is performed using the mean-field distribution in Eq. (46) at the scaled temperature $\tau=0.1$. As the SCE model, we consider the constant term and two-body interactions with the lowest $l$-index as
$\begin{split}&E_{\rm SCE}=\\ &J_{0}+\sum_{i\neq j}\sum_{l_{i},l_{j}=1}^{l_{\rm max}=1}\sum_{m_{i}=-l_{i}}^{l_{i}}\sum_{m_{j}=-l_{j}}^{l_{j}}J_{ij\bm{l}\bm{m}}Z_{l_{i}m_{i}}(\hat{\bm{e}}_{i})Z_{l_{j}m_{j}}(\hat{\bm{e}}_{j}).\end{split}$ (47)
@@
We included all possible two-body pairs (Fe–Fe, Fe–Ge, and Ge–Ge). By constructing symmetry-adapted basis functions, the number of independent SCE coefficients excluding the constant term resulted in 324. This SCE model is equivalent to the generalized Heisenberg model consisting of the isotropic exchange term, DMI, and anisotropic symmetric interaction term as
$\begin{split}E=&-\sum_{i,j,i\neq j}J_{ij}\hat{\bm{e}}_{i}\cdot\hat{\bm{e}}_{j}\\ &+\sum_{i,j,i\neq j}{\bm{D}}_{ij}\cdot(\hat{\bm{e}}_{i}\times\hat{\bm{e}}_{j})\\ &+\sum_{i,j,i\neq j}{\hat{\bm{e}}_{i}}^{\mathsf{T}}{\bm{\Gamma}}_{ij}\hat{\bm{e}}_{j},\end{split}$ (48)
@@
where $J_{ij}$ is the isotropic exchange parameter, ${\bm{D}}_{ij}$ is the DMI vector, and ${\bm{\Gamma}}_{ij}$ is the anisotropic symmetric interaction matrix with conditions ${\bm{\Gamma}}_{ij}={\bm{\Gamma}}_{ij}^{\mathsf{T}}$ and tr[$\tr[{\bm{\Gamma}}_{ij}]=0$. The conversion rule from SCE coefficients to $J_{ij}$, ${\bm{D}}_{ij}$, and ${\bm{\Gamma}}_{ij}$ is explained in Sec. IV.2.
Figure 2: Comparison of the convergence with respect to the number of DFT data points $N_{\mathrm{data}}$ for energy-fitted (black) and torque-fitted (red) regression. Panels (a) and (b) show the first-nearest-neighbor Fe–Fe exchange coupling $J_{01}$ and the magnitude of the DMI vector $|{\bm{D}}_{01}|$, respectively. The insets enlarge the torque-fitted regression results for $N_{\mathrm{data}}\leq 100$.
We first demonstrate that torque-fitted regression reduces the number of required DFT data points. Figure 2 plots the convergence of the first-nearest-neighbor Fe–Fe exchange coupling $J_{01}$ and the magnitude of the DMI vector $|\bm{D}_{01}|$ versus the number of DFT data points for the torque-fitted and energy-fitted approaches. In the conventional energy-fitted regression, both parameters begin to stabilize around $N_{\rm data}=\!210$, which is the number of independent parameters in the SCE model. In contrast, the torque-fitted regression requires far fewer DFT data points for parameter convergence, because each configuration provides $3N_{\rm a}$ targets (the $x$, $y$, and $z$ torque components on each atom) rather than a single scalar energy; thus, by a simple calculation, the number of configurations needed scales as $\sim 1/(3N_{\rm a})$ relative to the energy-fitted case—$1/192$ in the present system ($N_{\rm a}=64$). Consequently, convergence is achieved with only a small number of data points. Collectively, these reductions make the SCE model practical and data-efficient when parameterized via torque-fitted regression. As the supercell size grows (and thus $N_{\rm a}$ increases), the relative efficiency over the energy-fitted approach becomes even more pronounced, making torque-fitted parameterization especially advantageous for larger supercells.
Figure 3: (a) Spin stiffness constant AA, (b) spiralization constant $D$, and (c) helical spin period $\lambda$ in ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$. Results of previous theoretical works [45, 47, 44, 46] and experimental works [37, 42, 41, 43] are also presented for reference.
@@
Next, we analyze the helical spin structures of ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$. Following Refs. [45, 46, 47, 48], we employ a micromagnetic energy functional derived from Eq. (48) together with the B20 crystal geometry. Exploiting the cubic symmetry of the B20 structure, the energy functional can be written as [48]
@@
$A$ $=\cfrac{1}{2}\sum_{ij}J_{ij}{\bm{R}}_{ij}\cdot{\bm{R}}_{ij},$ (50)
$D$ $=\sum_{ij}{\bm{D}}_{ij}\cdot{\bm{R}}_{ij},$ (51)
@@
$\lambda=4\pi\left|\cfrac{A}{D}\right|.$ (52)
@@
Figures 3(a)–(c) show the composition dependence of A, D, and $\lambda$. Here, we emulated changes in chemical composition by electron and hole doping, and the lattice constant was fixed to that of $\rm FeGe$. The number of data $N_{\rm data}$ is set to 100. For the spin stiffness constant $A$ [Fig. 3(a)], we find sizable deviations from previous theoretical studies [45, 47] across the composition range of ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$. Given that the experimental magnetic transition temperatures of $\rm FeGe$ and $\rm MnGe$ are $\sim 280$ K and $\sim 170$ K, respectively, and that within mean-field theory critical temperature is proportional to $A$, the results of Grytsiuk et al. [47] capture the expected trend (they evaluated spin-model parameters by a perturbative approach), whereas our data exhibit the opposite tendency. A likely origin is the limited simulation cell used here ($\times2$$\times$2). Indeed, Ref. [12] reports that the nearest-neighbor $J_{ij}$ between Mn atoms in MnGe is roughly twice as large as the nearest-neighbor Fe–Fe $J_{ij}$ in $\rm FeGe$; nevertheless, the smaller transition temperature of MnGe can be understood as arising from contributions of longer-range $J_{ij}$. However, within the SCE framework this discrepancy can be systematically reduced by increasing the supercell size; the remaining choice becomes a trade-off between computational cost and accuracy.
@@
Regarding the parameter $D$ [Fig. 3(b)], its composition dependence agrees well with previous theoretical studies. In particular, we find the sign change of $D$ around $x(\mathrm{Fe})\approx 0.6$$0.7$ and $y(\mathrm{Co})\approx 0.6$. As noted above, the helical period $\lambda$ diverges at compositions where $D=0$. Figure 3(c) compares $\lambda$ calculated from $A$ and $D$ as Eq. (52) with experimental results. Experiments likewise show a divergence near $x(\mathrm{Fe})\approx 0.8$ and $y(\mathrm{Co})\approx 0.5$; aside from modest offsets, the overall trends are consistent. These offsets are reasonable in light of our computational setup—varying the electron number only while fixing the lattice constant—and fall within the expected uncertainty. Overall, although finite-size effects associated with the supercell used here are present, our calculations clearly indicate that the SCE framework attains systematically improved accuracy as the cell size is increased. In the following, we evaluate the impact of spin fluctuations—a representative temperature effect—where the advantages of SCE become particularly pronounced.
@@
Whether a spin model can accurately capture the impact of thermal magnetic disorder is a key requirement for quantitative modeling. In this regard, the SCE approach offers a complementary advantage to widely used perturbative schemes. Whereas perturbative methods are formulated around infinitesimal spin rotations about a chosen reference state, SCE constructs the spin model directly from datasets of genuinely disordered magnetic configurations. Following the $\tau=0.1$ results presented above, we increase the reduced temperature to $\tau=0.3$, sample 100 spin configurations, and construct the corresponding SCE model. We also construct an SCE model with the expansion cutoff in Eq. (47) raised from $l_{\max}=1$ to $l_{\max}=2$ as
$\begin{split}&E_{\rm SCE}=\\ &J_{0}+\sum_{i\neq j}\sum_{l_{i},l_{j}=1}^{l_{\rm max}=2}\sum_{m_{i}=-l_{i}}^{l_{i}}\sum_{m_{j}=-l_{j}}^{l_{j}}J_{ij\bm{l}\bm{m}}Z_{l_{i}m_{i}}(\hat{\bm{e}}_{i})Z_{l_{j}m_{j}}(\hat{\bm{e}}_{j}),\end{split}$ (53)
@@
Figure 4: (a) Parity plot of site-resolved torque components. $x=T_{i\alpha}^{\rm DFT}$, $y=T_{i\alpha}^{\rm SCE}$, $\alpha\in\{x,y,z\}$. (b) Histogram of site-resolved magnetic-moment magnitudes of Fe $|{\bm{M}}_{\rm Fe}|$ at $\tau=0.1$ (blue) and $\tau=0.3$ (orange); bin width $0.01\ \mu_{\rm B}$. (c) $\langle\theta_{i}\rangle$ defined in Eq. (55) plotted against Fe moment magnitude $|\bm{M}_{\rm Fe}|$. (d)(e) Parity plots of Fe torques at $\tau=0.3$ for $l_{\rm max}=1$ and $l_{\rm max}=2$, respectively. Points are color-coded by the deviation from the modal Fe moment magnitude, $|\bm{M}_{\rm Fe}|-|\bm{M}_{\rm Fe}|_{\rm mode}$. The modal value is $|\bm{M}_{\rm Fe}|_{\rm mode}=1.265\ \mu_{\rm B}$.
The blue and gray points in Figure 4(a) compare DFT torques with those from the $l_{\max}=1$ SCE model at $\tau=0.1$ and $\tau=0.3$, respectively, with all other settings identical, thereby isolating the effect of increased magnetic disorder. At $\tau=0.1$, the $l_{\max}=1$ SCE reproduces the DFT torques closely, whereas its predictive accuracy deteriorates at $\tau=0.3$. These trends indicate that truncation at $l_{\max}=1$ is valid at very low temperature range but insufficient at elevated temperature. By contrast, at $\tau=0.3$, extending the SCE to $l_{\max}=2$ substantially improves the torque prediction accuracy relative to $l_{\max}=1$. We attribute the reduced $l_{\max}=1$ accuracy at $\tau=0.3$ to the larger spread in the magnetic moment magnitude of Fe, $|{\bm{M}}_{\rm Fe}|$, as discussed below.
Figure 4(b) shows histograms of $|\bm{M}_{\rm{Fe}}|$ at $\tau=0.1$ and $\tau=0.3$ (Ge moments are negligible compared with Fe and are omitted). The $\tau=0.3$ distribution is visibly broader than at $\tau=0.1$, indicating larger dispersion in $|\bm{M}_{\rm{Fe}}|$. This broad variance in $|\bm{M}_{\rm{Fe}}|$ is strongly correlated with local magnetic disordering. To quantify short-range order among Fe–Fe pairs, we define for each Fe site $i$ the average relative angle to its first-nearest Fe neighbors,
@@
where ${\rm 1NN}(i)$ is the set of first-nearest Fe neighbors of site $i$ and $N_{\rm 1NN}(i)$ is the number of elements in the set ${\rm 1NN}(i)$. Figure 4(c) clearly shows that larger $\langle\theta_{i}\rangle$ leads to the shrinkage of $|\bm{M}_{\rm{Fe}}|$, i.e., $|\bm{M}_{\rm{Fe}}|$ strongly depends on local magnetic order. Figure 4(d) presents the $\tau=0.3$, $l_{\max}=1$ parity plot with points color-coded by the deviation from the modal Fe moment magnitude at $\tau=0.3$, $|\bm{M}_{\mathrm{Fe}}|-|\bm{M}_{\mathrm{Fe}}|_{\mathrm{mode}}$. Taken together, these results indicate that the deterioration of the $l_{\max}=1$ SCE torque predictions with increasing $\tau$ stems from its inability to capture the temperature-dependent spread in local-moment magnitudes $|{\bm{M}}_{\rm Fe}|$ induced by magnetic disordering. Moreover, the improvement of torque predictive accuracy by including higher-order interactions can be understood intuitively from Fig. 4(c); these terms provide the nonlinear angular dependence needed to reproduce the energy and torque for configurations with large relative spin angles behavior that cannot be captured by the low-order $l_{\max}=1$ model. Indeed, the torques on atoms with strongly shrunk local moments—poorly captured by the $l_{\max}=1$ model [Fig. 4(d)]—are reproduced much more accurately by the $l_{\max}=2$ SCE model [Fig. 4(e)].

### v2

@@ II.4 Real spherical harmonics represented as polynomials of Cartesian coordinates
In torque-fitting regression formalism, we must evaluate gradients of the spherical harmonics with respect to the unit vector, cf. Eq. (28). Differentiation in spherical coordinates $(\theta,\phi)$ is numerically ill-conditioned near the poles ($\sin\theta\!\to\!0$). To avoid these coordinate singularities, we represent the spherical harmonics as polynomials in the Cartesian components of the unit vector, which yields well-behaved analytic derivatives via $\partial/\partial e_{x}$, $\partial/\partial e_{y}$, and $\partial/\partial e_{z}$. The gradient of the complex spherical harmonics with respect to the unit vector was already derived in Ref. [31]. Here, we present expressions for the gradients of the real spherical harmonics in their Cartesian-polynomial form. See also the Supplemental Material for details of the derivations [32].
@@
III Application
@@
B20-type chiral magnetic compounds ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$ exhibit helical spin structures [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. These compounds have a moderate SOC strength, making them suitable testbeds for developing DFT-based methods to evaluate the DMI  [44, 45, 46, 47]. Here, we apply the SCE method to the B20-type compounds and examine the applicability via the evaluation of exchange interactions, DMI, and helical spin period.
III.1 Computational details
[…]
III.2 Sampling method based on the mean-field approximation
Constructing an SCE model requires sampling spin configurations that adequately span the magnetic states of interest. In general, building a model that reproduces energies across arbitrary magnetic states with high accuracy within a realistic computational budget will demand sophisticated strategies, which we defer to future work. In the applications presented below, we instead assume that the relevant magnetic states are known a priori. For this setting, we employ a practical mean-field-based sampling scheme for the Heisenberg model that efficiently generates spin configurations representative of the target states [12, 52, 53]. Within the mean-field Heisenberg model, the single-site orientational probability density $P(\hat{\bm{e}}_{i})$ takes the von Mises–Fisher form,
$P(\hat{\bm{e}}_{i})=\cfrac{3{\mathrm{m}}/\tau}{4\pi\sinh(3{\mathrm{m}}/\tau)}\exp\left(\cfrac{3{\mathbf{m}}\cdot\hat{\bm{e}}_{i}}{\tau}\right),$ (45)
where $^{\mathbf{m}}=\langle\hat{\bm{e}}\rangle={\mathrm{m}}\hat{\bm{e}}$ ($0\leq{\mathrm{m}}\leq 1$) is the site-independent thermal average of the spin and $\tau=T/T_{\rm c}^{\rm MFA}$ is a reduced temperature ($T$ is the temperature and $T_{\rm c}^{\rm MFA}$ is the mean-field critical temperature). The parameters $\mathrm{m}$} and $\tau$ are self-consistently related; fixing one automatically determines the other. Therefore, by choosing $\tau$, one can sample $\hat{\bm{e}}_{i}$ from this distribution to generate spin configurations representative of the target magnetic disorder within the mean-field approximation. Please refer to Ref. [12] for the derivation and further details. In this study, we used $\tau=0.1$ for demonstrating data efficiency (Sec. III.4) and for deriving micromagnetic parameters (Sec. III.5), while $\tau=0.3$ was additionally employed to evaluate higher-order interaction terms (Sec. III.6).
III.3 Spin-cluster expansion and micromagnetic models
For the SCE model, we first consider the constant term and two-body interactions with the lowest $l$-index as
$\begin{split}E_{\rm SCE}=J_{0}+\sum_{i,j}\sum_{l_{i},l_{j}=1}^{l_{\rm max}=1}\sum_{m_{i}=-l_{i}}^{l_{i}}\sum_{m_{j}=-l_{j}}^{l_{j}}J_{ij\bm{l}\bm{m}}Z_{l_{i}m_{i}}(\hat{\bm{e}}_{i})Z_{l_{j}m_{j}}(\hat{\bm{e}}_{j}).\end{split}$ (46)
@@ III.3 Spin-cluster expansion and micromagnetic models
The sum $\sum_{i,j}$ runs over all ordered pairs with $i\neq j$ (self-terms are excluded), and each pair is counted twice. We included all possible two-body pairs (Fe–Fe, Fe–Ge, and Ge–Ge). While the Ge moment is largely induced and therefore small, including Ge-related pairs avoids imposing an a priori truncation and lets their contributions be assessed quantitatively. By constructing symmetry-adapted basis functions, the number of independent SCE coefficients excluding the constant term resulted in 196. This SCE model is equivalent to the generalized Heisenberg model consisting of the isotropic exchange term, DMI, and anisotropic symmetric interaction term as
$\begin{split}E=&-\sum_{i,j}J_{ij}\hat{\bm{e}}_{i}\cdot\hat{\bm{e}}_{j}\\ &+\sum_{i,j}{\bm{D}}_{ij}\cdot(\hat{\bm{e}}_{i}\times\hat{\bm{e}}_{j})\\ &+\sum_{i,j}{\hat{\bm{e}}_{i}}^{\mathsf{T}}{\bm{\Gamma}}_{ij}\hat{\bm{e}}_{j},\end{split}$ (47)
@@
where $J_{ij}$ is the isotropic exchange parameter, ${\bm{D}}_{ij}$ is the DMI vector, and ${\bm{\Gamma}}_{ij}$ is the anisotropic symmetric interaction matrix with conditions ${\bm{\Gamma}}_{ij}={\bm{\Gamma}}_{ij}^{\mathsf{T}}$ and tr[$\tr[{\bm{\Gamma}}_{ij}]=0$. The conversion rule from SCE coefficients to $J_{ij}$, ${\bm{D}}_{ij}$, and ${\bm{\Gamma}}_{ij}$ is explained in the Appendix.
@@
Following Refs. [45, 46, 47, 54], we employ a micromagnetic energy functional derived from Eq. (47) together with the B20 crystal geometry. Exploiting the cubic symmetry of the B20 structure, the energy functional can be written as [54]
@@
$A$ $=\cfrac{1}{2}\sum_{i,j}J_{ij}{\bm{R}}_{ij}\cdot{\bm{R}}_{ij},$ (49)
$D$ $=\sum_{i,j}{\bm{D}}_{ij}\cdot{\bm{R}}_{ij},$ (50)
@@
$\lambda=4\pi\left|A/D\right|.$ (51)
@@
III.4 Data efficiency of the regression scheme
Figure 2: Comparison of the convergence with respect to the number of DFT data points $N_{\mathrm{data}}$ for energy-fitting (black) and torque-fitting (red) regression. Panels (a) and (b) show the first-nearest-neighbor Fe–Fe exchange coupling $J_{01}$ and the magnitude of the DMI vector $|{\bm{D}}_{01}|$, respectively. The insets enlarge the torque-fitting regression results for $N_{\mathrm{data}}\leq 100$.
We first demonstrate that torque-fitting regression reduces the number of required DFT data points. Figure 2 plots the convergence of the first-nearest-neighbor Fe–Fe exchange coupling $J_{01}$ and the magnitude of the DMI vector $|\bm{D}_{01}|$ versus the number of DFT data points for the torque-fitting and energy-fitting approaches. In the conventional energy-fitting regression, both parameters begin to stabilize around $N_{\rm data}=\!196$, which is the number of independent parameters in the SCE model. In contrast, the torque-fitting regression requires far fewer DFT data points for parameter convergence, because each configuration supplies $3N_{\rm a}$ torque components (the $x$, $y$, and $z$ components on each atom) rather than a single scalar energy. A simple counting argument therefore suggests a potential reduction in the required number of configurations by a factor on the order of $3N_{\rm a}$ relative to energy fitting (i.e., $\sim 1/192$ for the present system with $N_{\rm a}=64$), although the practical gain depends on correlations among torque components, symmetry constraints, and the resulting conditioning of the design matrix. Consistently, we observe that convergence is achieved with only a small number of data points. Collectively, these reductions make the SCE model practical and data-efficient when parameterized via torque-fitting regression. As the supercell size grows (and thus $N_{\rm a}$ increases), the relative efficiency over the energy-fitting approach becomes even more pronounced, making torque-fitting parameterization especially advantageous for larger supercells.
III.5 Evaluation of micromagnetic parameters
Figure 3: (a) Spin stiffness constant AA, (b) spiralization constant $D$, and (c) helical spin period $\lambda$ in ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$. The two red dashed lines indicate the compositions at which $\lambda$ diverges in this study. Results of previous theoretical works [45, 47, 44, 46] and experimental works [37, 42, 41, 43] are also presented for reference.
@@ III.5 Evaluation of micromagnetic parameters
Figures 3(a)–(c) show the composition dependence of $A$, $D$, and $\lambda$, respectively. The number of data $N_{\rm data}$ is set to 100. For the spin stiffness constant $A$ [Fig. 3(a)], we find sizable deviations from previous theoretical studies [45, 47] across the composition range of ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$. Given that the experimental magnetic transition temperatures of $\rm FeGe$ and $\rm MnGe$ are $\approx 280$ K [34] and $\approx 170$ K [35], respectively, and that within mean-field theory critical temperature is proportional to $A$, the results of Grytsiuk et al. [47] capture the expected trend (they evaluated spin-model parameters by a perturbative approach), whereas our data exhibit the opposite tendency. A likely origin is the limited simulation cell used here (the $\times2$$\times$2 supercell). Indeed, Ref. [13] reports that the nearest-neighbor $J_{ij}$ between Mn atoms in MnGe is roughly twice as large as the nearest-neighbor Fe–Fe $J_{ij}$ in $\rm FeGe$; nevertheless, the smaller transition temperature of MnGe can be understood as arising from contributions of longer-range $J_{ij}$. However, within the SCE framework this discrepancy can be systematically reduced by increasing the supercell size; the remaining choice becomes a trade-off between computational cost and accuracy.
@@
Regarding the parameter $D$ [Fig. 3(b)], its composition dependence agrees well with previous theoretical studies. In particular, we find the sign change of $D$ around $x(\mathrm{Fe})\approx 0.6$$0.7$ and $y(\mathrm{Co})\approx 0.6$. As noted above, the helical period $\lambda$ diverges at compositions where $D=0$. Figure 3(c) compares $\lambda$ calculated from $A$ and $D$ as Eq. (51) with experimental results. Experiments likewise show a divergence near $x(\mathrm{Fe})\approx 0.8$ and $y(\mathrm{Co})\approx 0.5$; aside from modest offsets, the overall trends are consistent. These offsets are reasonable in light of our computational setup—varying the electron number only while fixing the structural parameters—and fall within the expected uncertainty. Overall, although finite-size effects associated with the supercell used here are present, our calculations indicate that the SCE framework attains systematically improved accuracy as the cell size is increased. In the following, we evaluate the impact of spin fluctuations—a representative temperature effect—where the advantages of SCE become particularly pronounced.
III.6 Higher-order interactions for thermally disordered states
@@
Whether a spin model can accurately capture the impact of thermal magnetic disorder is a key requirement for quantitative modeling. Following the $\tau=0.1$ results presented above, we increase the reduced temperature to $\tau=0.3$, sample 100 spin configurations, and construct the corresponding SCE model. We also construct an SCE model with the expansion cutoff in Eq. (46) raised from $l_{\max}=1$ to $l_{\max}=2$ as
$\begin{split}E_{\rm SCE}=J_{0}+\sum_{i,j}\sum_{l_{i},l_{j}=1}^{l_{\rm max}=2}\sum_{m_{i}=-l_{i}}^{l_{i}}\sum_{m_{j}=-l_{j}}^{l_{j}}J_{ij\bm{l}\bm{m}}Z_{l_{i}m_{i}}(\hat{\bm{e}}_{i})Z_{l_{j}m_{j}}(\hat{\bm{e}}_{j}),\end{split}$ (52)
@@
Figure 4: Parity plot of site-resolved torque components: the horizontal axis shows $T_{{\rm DFT},i\alpha}$ and the vertical axis shows $T_{{\rm SCE},i\alpha}$, where $\alpha\in\{x,y,z\}$. (b) Histogram of Fe-moment magnitude $|\mathfrak{m}_{\rm Fe}|$ at $\tau=0.1$ (blue) and $\tau=0.3$ (orange); bin width $0.01\ \mu_{\rm B}$. (c) $|\mathfrak{m}_{\rm Fe}|$ as a function of $\langle\theta_{i}\rangle$ defined in Eq. (54). (d, e) Parity plots of Fe torques at $\tau=0.3$ for $l_{\rm max}=1$ and $l_{\rm max}=2$, respectively (same axis definitions as in panel (a)). In (d) and (e), points are color-coded by the deviation from the mode of Fe moment distribution, $|\mathfrak{m}_{\rm Fe}|-|\mathfrak{m}_{\rm Fe}|_{\rm mode}$with $|\mathfrak{m}_{\rm Fe}|_{\rm mode}=1.265\ \mu_{\rm B}$.
The blue and gray points in Figure 4(a) compare DFT torques with those from the $l_{\max}=1$ SCE model at $\tau=0.1$ and $\tau=0.3$, respectively, with all other settings identical, thereby isolating the effect of increased magnetic disorder. At $\tau=0.1$, the $l_{\max}=1$ SCE reproduces the DFT torques closely, whereas its predictive accuracy becomes less accurate at $\tau=0.3$. These trends indicate that truncation at $l_{\max}=1$ is valid at very low temperatures but insufficient at elevated temperatures. By contrast, at $\tau=0.3$, extending the SCE to $l_{\max}=2$ substantially improves the torque prediction accuracy relative to $l_{\max}=1$. We attribute the reduced $l_{\max}=1$ accuracy at $\tau=0.3$ to the larger spread in the magnetic moment magnitude of Fe, $|\mathfrak{m}_{\rm Fe}|$, as discussed below.
Figure 4(b) shows histograms of the Fe magnetic-moment magnitude, $|\mathfrak{m}_{\rm{Fe}}|$, at $\tau=0.1$ and $\tau=0.3$ (the Ge moments are negligible compared with those of Fe and are therefore omitted). The $\tau=0.3$ distribution is visibly broader than at $\tau=0.1$, indicating larger dispersion in $|\mathfrak{m}_{\rm{Fe}}|$. This broad variance in $|\mathfrak{m}_{\rm{Fe}}|$ is strongly correlated with local magnetic disordering. To quantify short-range order among Fe–Fe pairs, we define for each Fe site $i$ the average relative angle to its first-nearest Fe neighbors,
@@
where ${\rm 1NN}(i)$ is the set of first-nearest Fe neighbors of site $i$ and $N_{\rm 1NN}(i)$ is the number of Fe atoms in the set ${\rm 1NN}(i)$. Figure 4(c) clearly shows that larger $\langle\theta_{i}\rangle$ leads to the shrinkage of $|\mathfrak{m}_{\rm{Fe}}|$, i.e., $|\mathfrak{m}_{\rm{Fe}}|$ strongly depends on local magnetic order. Figure 4(d) presents the $\tau=0.3$, $l_{\max}=1$ parity plot with points color-coded by the deviation from the modal Fe moment magnitude at $\tau=0.3$, $|\mathfrak{m}_{\mathrm{Fe}}|-|\mathfrak{m}_{\mathrm{Fe}}|_{\mathrm{mode}}$. Taken together, these results indicate that the degradation of the $l_{\max}=1$ SCE torque predictions with increasing $\tau$ stems from its inability to capture the temperature-dependent spread in local-moment magnitudes $|\mathfrak{m}_{\rm Fe}|$ induced by magnetic disordering. Moreover, the improvement of torque predictive accuracy by including higher-order interactions can be understood intuitively from Fig. 4(c); these terms provide the nonlinear angular dependence needed to reproduce the energy and torque for configurations with large relative spin angles, which cannot be captured by the low-order $l_{\max}=1$ model. Indeed, the torques on atoms with strongly shrunk local moments—poorly captured by the $l_{\max}=1$ model [Fig. 4(d)]—are reproduced much more accurately by the $l_{\max}=2$ SCE model [Fig. 4(e)].